OpenStudy (anonymous):

two particles of masses 1g and 4g are moving with equal kinectic energy.what is the ratio of their linear momenta.

OpenStudy (anonymous):

here are your two equations 1~~~ (1/2)*(m1)(v1^2)=(1/2)*(m2)(v2^2) and 2~ (ratio of linear momentum) (m1*v1)/(m2*v2). IMPORTANT NOTE: you can invert this and get a ratio of 2 to 1 instead of 1 to 2, idk what way you want it.

OpenStudy (anonymous):

So you solve in terms of V1 persay, V1= sqrt(m2/m1)*V2 plugging this into your ratio gives : [m1*sqrt(m2/m1)*v2]/[m2*v2] where v2 cancels out giving [m1*sqrt(m2/m1)]/[m2]

OpenStudy (anonymous):

the answer is 1/2, so how does your equation reach 1/2?